[Hallc_running] [Revised Logentry] RC update

alis at jlab.org alis at jlab.org
Sat Dec 11 13:45:02 EST 2021

Logentry Text:
RC meeting at 10 AM
To join the Meeting:

-Beam ran well yesterday at roughly 5000 pions at end of owl. Finished with LD2 runs, moving on to HEEP singles.
-Adjusted SHMS HMS settings to prepare for heep singles, changed target to LH2
-Q3 of SHMS was down before day shift, 7:20 AM, 
-Amy Comer, magnet on-call expert, was called in to power cycle the magnet.
-Beam is ready for hall, magnet has ramped up successfully after reset
-Asked for beam to hall (9:45 AM) and ready to go
-Target pressure for loop 2, LH2, is lower than usual - around 17-18 psi. The alarm went off early in the day shift which brought it to the day shift target operator's attention (Bill Henry). Usually acceptable pressure is around 22 psi. If this drops to atmospheric pressure, 14.7 psi, then we have a problem. The pressure was around this range 17-18 psi for LD2 as well. We did contact target expert JP to resolve issue but it is currently ongoing and UNRESOLVED. We assume we can continue to take data. Note that we will be taking LH2 data at 70 uA today.

For details see


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3960802
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211211/5208f3a4/attachment.html>

More information about the Hallc_running mailing list